VariableContribution Fixed Costs Income (Loss) Units Sold Selling Price Operating Costs Per Unit Margin 10,900 S 21.00 7,700 Firm A Firm B Firm C Firm D S 105,000 42,200 66,900 32.330 (6,660) 18.00 7.20 460 10,500 4.60 4,930 54.00| 42,050| 49,810Calculate the missing amounts for each of the following firms: (Do not round intermediate calculations. Round "Selling Price" and "Variable Costs Per Unit" to 2 decimal places.)
Answer #1

* Requirement , as asked

Unit Sold Selling price Variable cost per unit Contribution margin Fixed Cost Operating Income (Loss)
Firm A                                              10,900 $                   21.00 $                 11.37 $            105,000 $               42,200 $              62,800
Firm B                                                 7,700 $                   30.89 $                 18.00 $               99,230 $               66,900 $              32,330
Firm C                                                 4,038 $                     7.20 $                    4.60 $               10,500 $               17,160 $              (6,660)
Firm D                                                 4,930 $                   62.53 $                 54.00 $               42,050 $               49,810 $              (7,760)

---Working

Unit Sold Selling price Variable cost per unit Contribution margin Fixed Cost Operating Income (Loss)
Firm A 10900 21 =21-(105000/10900) 105000 42200 =105000-42200
Firm B 7700 =+(99230/7700)+18 18 =66900+32330 66900 32330
Firm C =10500/(7.2-4.6) 7.2 4.6 10500 =10500-(-6660) -6660
Firm D 4930 =+(42050/4930)+54 54 42050 49810 =42050-49810